OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (anonymous):

kevsturge 0 (1,5) (-2,3) Difference in y/x = y= 5-3 = 2 x= 1-(-2) = 3 gradient = 2/3 y=2/3x+c substitute one of the points for x any y (-2,3) gives: 3=2/3(-2)+c ; rearrange to give: 3 = -4/3 + c so c is 3+4/3 = c 13/3 = c Now put all together to give the answer y=2/3x+13/3

OpenStudy (anonymous):

sorry had a funny 5 minutes: y=2/3x+13/3 is the answer